OpenStudy (anonymous):

What is the distance between points (-2, -4) and (4, -7)?

pooja195 (pooja195):

We use the distance formula \[\huge d=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\]

pooja195 (pooja195):

(x1,y1)(x2,y2)

OpenStudy (arindameducationusc):

https://www.studyblue.com/notes/note/n/math-formulas/deck/5285669 Use distance formula

OpenStudy (arindameducationusc):

wait.. i will upload something

pooja195 (pooja195):

@arindameducationusc i already wrote it

OpenStudy (arindameducationusc):

These steps... co- relate with the formula.....

OpenStudy (arindameducationusc):

@pooja195 Could you please help in my question?

OpenStudy (anonymous):

Use the distance formula, \[distance = \sqrt{(x _{2}-x_{1})^{2}+(y_{2}-y_{1})^{2}}\] plug in your coordinates, \[distance = \sqrt{(4--2)^{2}+(-7--4)^{2}}\] then solve the equation.